OpenStudy (megannicole51):

Find the integral of ln^5(z)/z

OpenStudy (anonymous):

make a u substitution \[u=\ln(z), du =\frac{dz}{z}\] giving you \[\int u^5du\]

OpenStudy (anonymous):

now i know you remember the power rule backwards ! \[\int u^5du=\frac{u^6}{6}\] right?!

OpenStudy (megannicole51):

yes i remember:D

OpenStudy (anonymous):

then replace \(u\) by \(\ln(z)\) and get \[\frac{1}{6}\ln^6(z)\]

OpenStudy (anonymous):

and dont forget the stupid \(+C\) out at the end...

OpenStudy (megannicole51):

thank you :D i just need to keep practicing these

OpenStudy (anonymous):

btw turning \(\int u^5du=\frac{u^6}{6}\) is the power rule backwards that u-substitution is the chain rule backwards

OpenStudy (anonymous):

\[(f(g(x))'=f'(g(x))g'(x)\] \[\int f'(g(x))g'(x)dx=f(g(x))\]

OpenStudy (anonymous):

of course you had to recognize \(\frac{1}{z}\) as the derivative of \(\ln(z)\)

OpenStudy (megannicole51):

i cant put ln into my online stuff so how would i be able to write ln^6 a different way?

OpenStudy (megannicole51):

(1/x)?

OpenStudy (anonymous):

i am not sure exactly what you are asking

OpenStudy (anonymous):

you have to put this in a computer homework? try (ln(z)^6)/6

OpenStudy (megannicole51):

my homework wont let me put in the ln^6 bc it doesnt recognize ln

OpenStudy (anonymous):

you can't write "ln" you have to write "ln(z)"

OpenStudy (anonymous):

like i said, try "(ln(z)^6)/6+C" that should work

OpenStudy (anonymous):

just like you can't write "sin" you have to write "sin(x)"

OpenStudy (megannicole51):

perfect:) thank you!

OpenStudy (megannicole51):

would i be able to solve a problem if there is a square root in the denominator?

OpenStudy (anonymous):

i bet it if is in the book or on line you can solve it you got a specific example?

OpenStudy (megannicole51):

its.....4e^4squareroot of y divided by square root y

OpenStudy (anonymous):

\[\int\frac{ 4e^{4\sqrt{x}}}{\sqrt{x}}dx\]?

OpenStudy (anonymous):

lets take that first 4 out front first and write \[4\int \frac{e^{4\sqrt{x}}}{\sqrt{x}}dx\]

OpenStudy (megannicole51):

okay:) easy enough lol

OpenStudy (anonymous):

another u-substitution \(e^{4\sqrt{x}}\) is a composite function the "outside function " being \(e^u\) and the "inside function" being \(4\sqrt{x}\)

OpenStudy (anonymous):

and it just so happens that the derivative of the inside function \(4\sqrt{x}\) is almost staring us right in the face the derivative of \(4\sqrt{x}\) is \(\frac{2}{\sqrt{x}}\)

OpenStudy (anonymous):

you don't have \(\frac{2dx}{\sqrt{x}}\) you have \(\frac{dx}{\sqrt{x}}\) no matter, it is just a constant, we can adjust

OpenStudy (anonymous):

ready ?

OpenStudy (anonymous):

or did i lose you?

OpenStudy (megannicole51):

how did u get 2/root x from 4*root x?

OpenStudy (megannicole51):

jk i know what u did

OpenStudy (anonymous):

i took the derivative of \(4\sqrt{x}\) and got \(\frac{2}{\sqrt{x}}\) ok? we need that, it is important

OpenStudy (megannicole51):

okay im good with that now:) proceed...

OpenStudy (anonymous):

ok here goes put \(u=4\sqrt{x}, du =\frac{2}{\sqrt{x}}dx\) and so \(\frac{1}{2}du=\frac{dx}{\sqrt{x}}\)

OpenStudy (anonymous):

making the substitution gives \[2\int e^udu\]

OpenStudy (anonymous):

the 2 because we divided the 4 out front by the 2 in \(\frac{1}{2}du\)

OpenStudy (anonymous):

now this is a rather easy integral right? since \(e^u\) is its own anti derivative and also its own anti derivative so \[2\int e^udu=2e^u\]

OpenStudy (megannicole51):

okay so far so good

OpenStudy (anonymous):

ok we are done when we go back and replace \(u\) by \(4\sqrt{x}\) and get a final answer of \[2e^{4\sqrt{x}}+C\]

OpenStudy (megannicole51):

wait how did we get rid of the radical in the denominator?

OpenStudy (anonymous):

your best hope of undersanding this gimmick is to take the derivative of \[2e^{4\sqrt{x}}\] and see that in fact you get \[\frac{4e^{4\sqrt{x}}}{\sqrt{x}}\]

OpenStudy (anonymous):

we got rid of the radical in the denominator when we wrote \[\frac{1}{2}du=\frac{dx}{\sqrt{x}}\]

OpenStudy (megannicole51):

the only part i don't understand is that

OpenStudy (megannicole51):

where did the 1/2 come from

OpenStudy (megannicole51):

actually i see it now

OpenStudy (anonymous):

here is what i would like you to do really, if you want to understand this trick take the derivative and see that you get what you want then you will understand the u substitution i can walk you through it if you like

OpenStudy (megannicole51):

sure:)

OpenStudy (anonymous):

ok lets ignore the annoying 4 out front and just take the derivative of \[e^{4\sqrt{x}}\]

OpenStudy (megannicole51):

oh and just a quick question....so in my notes theres a problem e^(3x)/7+e^(3x) and it says i cant do it but im not sure why

OpenStudy (anonymous):

by the chain rule, you get \[e^{4\sqrt{x}}\frac{d}{dx}[4\sqrt{x}]\] which is \[e^{4\sqrt{x}}\times \frac{2}{\sqrt{x}}\]

OpenStudy (megannicole51):

okay:) got it....keep join

OpenStudy (megannicole51):

*goin

OpenStudy (anonymous):

well what is another way to write this? it is the same as \[\frac{2e^{4\sqrt{x}}}{\sqrt{x}}\]

OpenStudy (megannicole51):

okay i see it

OpenStudy (anonymous):

so when you take the derivative of the composite function out pops the derivative of the inside function you are now trying to go backwards

OpenStudy (anonymous):

so instead of it popping out, it pops in !

OpenStudy (anonymous):

of course this trick only works if you are looking at \[f'(g(x))g'(x)\]

OpenStudy (anonymous):

in other words, a composite function with the derivative of the inside function staring you in the face somewhere

OpenStudy (megannicole51):

that actually makes sense

OpenStudy (anonymous):

imagine!

OpenStudy (anonymous):

all this garbage is just tricks and gimmicks to find a function whose derivative is given largely a waste of time, but you have to get through it most of what you need is written as formulas on the back page of your text

OpenStudy (megannicole51):

yeah and i wish professors would teach you the easy tricks instead of making us suffer

OpenStudy (anonymous):

math = suffer ok not really math + math teacher = suffer math = fun

OpenStudy (anonymous):

hey, at least it is on line and you can cheat right?

OpenStudy (megannicole51):

omg i completely forgot about wolfram!

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

now we will never see you here again will we?

OpenStudy (anonymous):

could have saved your self like an hour at least but you did learn something i hope

OpenStudy (megannicole51):

nah i will have a lot more questions bc i like people showing me step by step....if i just had someone show me an example and just write out the steps usually i could figure it out

OpenStudy (anonymous):

k see you later then good luck!

OpenStudy (megannicole51):

thank you so much for all of your help!